Juan is considering two independent projects. Project A costs $50,000 and cash flows of $18,700, $46,300, and $12,200 for Years 1 to 3, respectively. Project B costs $70,000 and has cash flows of $10,600, $15,800, and $67,900 for Years 1 to 3, respectively. Project C costs $25,000 and has cash flows of $15,300, $3,000, and $20,000 for Years 1 to 3, respectively. Juan assigns a discount rate of 10 percent to these projects. The total budget is $120,000. Which combinations of rojects is the best?
A. A+B
B. A+C
C. B+C
D. A+B+C
Answer #1

fc =SUM(D3:06) Year 1 Project A 2 Year Cash flows PV@10% 0 $ -50,000.00 1 $ 18,700.00 2 $ 46,300.00 3 $ 12,200.00 NPV Rank 1fx =SUM(D3:06) - D E G H L MN А 1 Project A F Project B Year K Project B Year 30 4 1 Cash flows Pv@10% -50000 18700 =C3/1.1 4